OpenStudy (anonymous):

Crystal is organizing the seating arrangements for a concert in her local park. The first row's seats are modeled by the function f(x) = 12. To give everyone a good view, each row must have three more seats than the row before it. The number of additional seats a row has is modeled by the function s(x) = 3(x - 1). Explain to Crystal how to create an equation to predict the number of seats in any row. Describe how to determine the number of seats in the 30th row. Use complete sentences

OpenStudy (anonymous):

@mathmate @ganeshie8 @Loser66 @e.mccormick @eyust707 @Preetha @wolfe8 @iPwnBunnies @paki @ShadowLegendX @Joseph91 @Hero @ikatouni @ByteMe @Loujoelou @ltp1212 @paki @neer2890 @Passionateboyy14 @perl @Mathfinder2014 @mathstudent55

OpenStudy (perl):

so lets call g(x) the seats in row x. g(x) = 12 + 3(x-1)

OpenStudy (perl):

so lets call g(x) the formula to find the number of seats in row x. g(x) = 12 + 3(x-1) THen find g(30) = 12 + 3*(30-1) = 99